OpenStudy (anonymous):

need help solving with elimination method. 8s+12t=16 6s+9t=12

hero (hero):

Hint: Simplify both equations first.

hero (hero):

Notice that if we divide the first equation by 4 and the second equation by 3, we can simplify both equations to this: 2s + 3t = 4 2s + 3t = 4

OpenStudy (anonymous):

m horrible with math I have no idea how and im so confused

hero (hero):

I understand.

OpenStudy (anonymous):

thank you so much for your help

hero (hero):

So if we tried to use the elimination method for this particular problem, we would end up with 0 = 0. Anytime you end up with that kind of result, it means that the solution to the system is all points (s,t) on either line.

hero (hero):

If that confuses you, just remember that when we simplified both equations, we ended up with: 2s + 3t = 4 2s + 3t = 4 which essentially means that both equations represent the same line. Therefore they both must have the same solutions.

OpenStudy (anonymous):

ok It makes a little sense on how you put it thank you

OpenStudy (anonymous):

I have one more if I post it do you mind helping me understand it? its using the same method ;)

hero (hero):

Well, unfortunately, not every system is simplifyable, and not every system will end up being representations of the same line.

hero (hero):

Such cases only represent a particular class of systems of equations.

hero (hero):

Namely "equivalent" equations.

hero (hero):

You will either have one of the three things: 1. a NORMAL system (two lines that intersect at a single point) 2. an INCONSISTENT system (two parallel lines that never intersect) 3. an EQUIVALENT system (two equations that represent the same line)

hero (hero):

So you will have different approaches to each.

OpenStudy (anonymous):

I just have problems trying to figure out what numbers to use. in the equations. now what do i do if they are not easy to eliminate like this one?

hero (hero):

Show me one you think is not easy. As far as I'm concerned, according to your situation, none of them are easy.

OpenStudy (anonymous):

ok let me post it in the question part so you get the points for helping me with it

hero (hero):

Okay, but you should close the question immediately afterwards.